OpenStudy (anonymous):

What is 0.2% of 20? I got 0.4 but the internet says its 0.04

OpenStudy (anonymous):

0.2% = 0.002 0.002 * 20 =

OpenStudy (anonymous):

0.002 * 20 = 0.02 * 2 = ...

OpenStudy (anonymous):

how exactly do you get from 0.2 to 0.002?

OpenStudy (anonymous):

\[ 0.2\% = \frac {0.2}{100} = \frac{0.02}{10} = \frac{0.002}{1} = 0.002 \]

OpenStudy (anonymous):

ok, I see. Why do we put 10 under 0.02?

OpenStudy (anonymous):

0's transfer over one at a time.

OpenStudy (anonymous):

ok. I see.